A sample of gas occupies 175mL at 10.0C. What volume in mL will the gas have at 20.0C and 750mm Hg?


1
Expert's answer
2021-05-19T02:26:57-0400

According to the gas law:

p1V1/T1 = p2V2/T2

where p1 - initial pressure, V1 - initial volume, T1 - initial temperature (10.0 °C = 283.15 K), p2 - final pressure, V2 - final volume, T2 - final temperature (20.0 °C = 293.15 K).

If the pressure is not changed:

V1/T1 = V2/T2

From here:

V2 = V1T2/T1 = 175 mL × 293.15 K / 283.15 K = 181.2 mL


Answer: 181.2 mL